# Communication Efficient LLM Pre-training with SparseLoCo

### v1

@@ 4.2 SparseLoCo
Table 2 compares SparseLoCo at against existing methods. We utilize 2-bit quantization for SparseLoCo with no observed loss degradation, while using the prescribed quantization settings for baselines Douillard et al. (2023a); Peng et al. (2024). We observe that SparseLoCo obtains lower final loss than DiLoCo and DeMo baselines, while enjoying the simultaneous communication benefits of aggressively sparsified pseudo-gradients and reduced synchronization frequency. As SparseLoCo inherently utilizes error feedback, SparseLoCo further reduces communication size by quantizing the sparsified values. We further compare the performance on simple downstream tasks relevant at this model scale in Table 3, demonstrating that the performance improvements are consistent.
In Figure 1, we further demonstrate the performance improvements of SparseLoCo at increasing values compared to well-tuned DiLoCo and DiLoCo without outer momentum baselines. We observe that SparseLoCo can consistently obtain a better loss than DiLoCo in each case. Finally, following Charles et al. (2025), we study an overtraining regime with a doubled token budget and a large communication interval (), where SparseLoCo still outperforms DiLoCo (see Table 8 of the Appendix).
SparseLoCo’s performance at different communication intervals (H)
@@
In Figure 1, we observe a trend that aligns with the hypothesis that SparseLoCo’s Outer EF can provide similar benefits as a DiLoCo outer momentum. In particular, we first observe that not using outer momentum in DiLoCo leads to significant performance degradation and that this setting corresponds exactly to the fully dense case for . With SparseLoCo, we observe that (i) extreme sparsity levels (when nearly nothing is sent, e.g., 0.05%) degrade performance. (ii) with increasing density (while remaining sparse), performance improves and eventually exceeds DiLoCo for all values of . (iii) finally, as approaches dense communication, the EF buffer becomes more sparse (due to line 12 in Algorithm 1), trending towards the performance of DiLoCo with no outer momentum and again degrading performance.
SparseLoCo can outperform DiLoCo and DiLoCo-LOM
@@
Across all settings of the inner steps (), we observe regimes where SparseLoCo outperforms both DiLoCo and DiLoCo-LOM. A plausible explanation is that sparse aggregation at a well-chosen emphasizes high-saliency components and reduces interference among updates, echoing intuitions from recent model-merging work in multi-task fine-tuning (Yadav et al., 2023; Davari & Belilovsky, 2024).
Higher sparsity is needed with fewer inner steps
@@
We observe through Figure 1 that the optimal value is reached at a higher sparsity level with fewer inner steps. This is consistent with the fact that higher inner steps communicate information from a larger total number of samples. Indeed, we would expect that a trajectory with more steps would have a larger support.
SparseLoCo is at the Pareto frontier in communication volume
@@
In Figure 2 we compare the communication volume of SparseLoCo to DiLoCo, DiLoCo w.o. outer momentum, and DeMo. The exact communication setting and the underlying implementation of the aggregation can have a significant impact on the communication volume. We consider two common setups from the literature—methods utilize either ring all-reduce or ring all-gather (Fig. A), or a parameter server (Fig. B). We observe that in both cases, SparseLoCo lies on the Pareto frontier while other methods have a strictly worse trade-off. We note that the results in Fig. A assume aggregation using a naive all-gather operation for implementation, while there is further potential to exploit the structure of the problem, for example, by summing overlapping indices along steps in the all-gather ring or utilizing specially designed all-reduce (Li & Hoefler, 2022). In Section A of the Appendix, we also discuss the communication measured during a live deployment of collaborative learning over the internet using SparseLoCo.
@@
Although our analysis and motivation in the work focuses on aggressively compressing the per-iteration message size we note that in communication settings where efficient all-reduce is already available and preferred, SparseLoCo still provides significant benefit over DiLoCo while incurring no additional memory or compute overhead. Concretely, the aggregation step in Algorithm 1 Line 14 can be performed directly by an all-reduce over a sparse vector. This has two significant benefits over DiLoCo with AR: (1) As observed in Table 8 and Figure 1, the performance when is optimally selected is improved over DiLoCo (2) the Outer error feedback naturally supports more aggressive quantization than the naive DiLoCo, allowing for 2-bit quantization to be used without an additional accumulator, unlike Thérien et al. (2025).
| Method | Density | Loss | Pseudo-Grad Size | # of Syncs | Quantization |
| AdamW DDP | 100% | 2.70 | 1.03 GB | 2445 | 16-bit |
| DiLoCo (H=15) | 100% | 2.76 | 0.48 GB | 163 | 8-bit |
| DeMo | 0.78% | 2.83 | 8.5 MB | 2445 | 8-bit |
| DeMo | 3.12% | 2.86 | 32.5 MB | 2445 | 8-bit |
| SparseLoCo (H=15) | 0.78% | 2.80 | 5.5 MB | 163 | 2-bit |
| SparseLoCo (H=15) | 3.12% | 2.73 | 17 MB | 163 | 2-bit |
| Method | arc_easy | hellaswag | piqa |
| AdamW DDP | 43.31% | 35.95% | 65.56% |
| DiLoCo | 44.95% | 34.14% | 65.07% |
@@
| SparseLoCo | 45.54% | 35.05% | 66.00% |
@@ 4.3 Ablations
A natural way to combine DiLoCo with OuterEF is by adding an error feedback as well as a Nesterov outer optimizer, as has been attempted by Thérien et al. (2025), who showed it can provide benefits for quantization but not sparsification. This approach requires an additional accumulator. In contrast, our finding is that, in the case of a high sparsification, using a global outer momentum can be detrimental to performance. This is illustrated in Table 4, where we equip SparseLoCo’s outer optimizer with Nesterov outer momentum. This significantly degrades performance at high sparsity. We hypothesize that this is due to the conflicting directions of the error feedback momentum and the outer momentum, since the largest components become amplified by the outer momentum but not the error feedback.
| Method | Loss |
| SparseLoCo | 2.72 |
| SparseLoCo + Nesterov Outer optimizer | 3.39 |
@@
We ablate the choice of compared to the alternative Random- (Shi et al., 2019; Wang et al., 2023) in Table 5. We observe that performance is significantly degraded when using random- for the same number of indices selected, emphasizing the importance of this design choice.
| Density | Random- Loss | Loss |
| 1.56% | 3.05 | 2.75 |
| 3.12% | 2.98 | 2.72 |
| 6.25% | 2.93 | 2.75 |
@@
As discussed, SparseLoCo benefits from stronger quantization than non-EF methods and supports up to 2-bit quantization and was generally observed to give results very close to full precision. In Table 6 below, we show the performance at different quantization values, showing that 2-bit quantization can be achieved at almost no performance cost.
| Quantization bits | Loss |
| 1 | 4.98 |
| 2 | 2.73 |
| 3 | 2.73 |
| 4 | 2.73 |
| 32 | 2.72 |
@@ 5 Conclusion
In conclusion, our work establishes that the outer momentum in DiLoCo can be replaced by local momentum accumulators without losing performance. Connecting local momentums with error feedback, we leverage this insight to develop SparseLoCo. Our algorithm can blend multi-iteration LLM pre-training methods with sparsification and quantization, enabling aggressive compression of DiLoCo’s pseudo-gradients. Our extensive experiments confirm that SparseLoCo significantly reduces communication while outperforming strong baselines such as DiLoCo and DeMo, placing it on the Pareto frontier of loss versus communication volume. Additionally, our experiments—outperforming DiLoCo—reveal that sparse aggregation may actually be crucial for improving the performance of multi-iteration, multi-worker methods compared to standard data-parallel pre-training. These results highlight the practical impact of SparseLoCo for communication-efficient LLM pre-training and suggest a path forward to scalable, low-bandwidth pre-training across data centers and globally distributed networks.
@@ Appendix A Real-World Deployment for Collaborative Permissionless Distributed Training over the Internet
SparseLoCo has been deployed in a real-world setting and used to collaboratively train models up to 8B with permissionless global participants using an incentive scheme Lidin et al. (2025) that rewards participants purely based on analysis of their compressed pseudo-gradients. This was done through the Bittensor blockchain and the templar project. The addition and coordination of peers and their rewards being handled through the blockchain. Communication of pseudo-gradients was routed through globally distributed, S3-compliant object storage—specifically Cloudflare R2—which enabled rapid dissemination of model updates worldwide. This setup allowed updates to be time-stamped and verified as part of the reward mechanism Lidin et al. (2025). Each peer maintained their own storage bucket, posting read credentials to a blockchain so that both other peers and the reward mechanism could access their compressed pseudo-gradients.
The deployment used replicas each associated to a peer, inner steps and a global batch size of 8M tokens per inner step. Although the system design allows peers to use any target hardware that can achieve reasonable throughput, the suggested hardware requirements targeted an H200.
SparseLoCo is particularly advantageous in this communication setup, as cloud providers have large bandwidth for peer downloads and are able to rapidly distribute and mirror files across the globe. For upload, a peer only sends their pseudo-gradient through the cloud provider. Therefore, their outbound communication (and required upload bandwidth) is kept low. They then download the pseudo-gradients from the cloud provider, which is able to easily handle the high bandwidth constraints. An example of a practical communication time measured with the 8B model is on average 12 seconds, including sending their compressed pseudo-gradients and downloading other workers’ messages with the test node never exceeding 500 Mb/s. Compared to processing with H200, which takes around 4.5 minutes, leading to minimal wall-clock time degradation despite traffic over the internet. For reference, Jaghouar et al. (2024), which trained a similarly sized model (10B) with 8-bit DiLoCo, reports a globally distributed all-reduce synchronization time of an average 8.3 minutes for a peak of nodes participating and processing time of 38 minutes. We also performed test measurements of communication time for a 70B model with the same setup as above ( peers), measuring a total communication time of 70 seconds on average, with the test node never reaching more than 500 Mb/s downlink and 110Mb/s uplink.
@@ Appendix B Effect of Chunking, DCT, and inner steps
A recently introduced method Peng et al. (2024) considered the single-step setting with error feedback, using a compression function that first applies a discrete cosine transform (DCT) on tensor chunks and then selects the values in the DCT domain. It further employed sign descent on the final aggregated update Kunstner et al. (2023). Without the DCT transform, this approach can be seen as a special case of SparseLoCo when , the inner optimizer is plain SGD, and the outer optimizer utilizes sign descent. Since the effect of the DCT transform, designed for data with sequential structure where the order of elements matters, is not well understood in this context, and given the additional uncertainty about the role of chunking, in this section we disentangle the contributions of both for DeMo, as well as in the multi-step () setting. The ablations of these three factors, evaluated purely in terms of loss, are presented in Table 7. Here, the EF baseline is a simplified DeMo that applies selection globally to the entire tensor (rather than within chunks) while still utilizing sign descent.
@@
Notably, DeMo does not have a natural way to incorporate adaptive optimization, and in practice, the sign descent is used to approximate the benefits of the Adam optimizer Peng et al. (2024). A significant advantage of SparseLoCo is that operating on the pseudo-gradients allows easy integration of adaptive optimizers like Adam in the inner loop.
@@ Appendix C Overtraining Regime with Large Communication Interval
Following Charles et al. (2025), we put SparseLoCo to the test in an overtraining regime by doubling the token budget to 20B and using a larger communication interval of . Our observations are consistent with the trends in Figure 1, and SparseLoCo outperforms DiLoCo at this setting (Table 8).
@@
Appendix D Hyperparameter Selection Details
In Table 9, we show the hyperparameter search spaces. We tune all methods at and reuse the best settings for other ; for DiLoCo, we also sweep , and use its best setting for higher . General and model settings are fixed across all runs unless stated otherwise.

### v2

@@ 4.2 SparseLoCo
Table 2 compares SparseLoCo at against existing methods. We utilize 2-bit quantization for SparseLoCo with no observed loss degradation, while using the prescribed quantization settings for baselines (douillard2023diloco; peng2024decoupled). We observe that SparseLoCo obtains lower final loss than DiLoCo and DeMo baselines, while enjoying the simultaneous communication benefits of aggressively sparsified pseudo-gradients and reduced synchronization frequency. As SparseLoCo inherently utilizes error feedback, SparseLoCo further reduces communication size by quantizing the sparsified values. We further compare the performance on simple downstream tasks relevant at this model scale in Table 3, demonstrating that the performance improvements are consistent.
@@
SparseLoCo’s performance at different communication intervals (H) In Figure 1, we further demonstrate the performance improvements of SparseLoCo at increasing values compared to well-tuned DiLoCo and DiLoCo without outer momentum baselines. We observe a trend that aligns with the hypothesis that SparseLoCo’s OuterEF can provide similar benefits to DiLoCo’s outer momentum. In particular, we first observe that not using outer momentum in DiLoCo leads to significant performance degradation and that this setting corresponds exactly to the fully dense case for (e.g., ). With SparseLoCo, we observe that (i) extreme sparsity levels (when nearly nothing is sent, e.g., ) degrade performance. (ii) With increasing density (while remaining sparse), performance improves and eventually exceeds DiLoCo for all values of . (iii) Finally, as approaches dense communication, the EF buffer becomes more sparse (due to line 12 in Algorithm 1), trending towards the performance of DiLoCo with no outer momentum and again degrading performance.
@@
SparseLoCo can outperform DiLoCo Across all settings of the inner steps , we observe regimes where SparseLoCo outperforms DiLoCo. A plausible explanation is that sparse aggregation at a well-chosen emphasizes high-saliency components and reduces interference among updates, echoing intuitions from recent model-merging work in multi-task fine-tuning (yadav2023tiesmerging; davari2024model).
@@
Higher sparsity is needed with fewer inner steps We observe through Figure 1 a systematic pattern that the optimal value of SparseLoCo is reached at a higher sparsity level with fewer inner steps. This is consistent with the fact that higher inner steps communicate information from a larger total number of samples. Indeed, we would expect that a trajectory with more steps would have a larger support.
@@
SparseLoCo is at the Pareto frontier in communication volume In Figure 2 we compare the communication volume of SparseLoCo to DiLoCo, DiLoCo w.o. outer momentum, and DeMo. The exact communication setting and the underlying implementation of the aggregation can have a significant impact on the communication volume. We consider two common setups from the literature—methods utilize either ring all-reduce or ring all-gather (Fig. A), or a parameter server (Fig. B). We observe that in both cases, SparseLoCo lies on the Pareto frontier while other methods have a strictly worse trade-off. We note that the results in Fig. A assume aggregation using a naive all-gather operation for implementation, while there is further potential to exploit the structure of the problem, for example, by summing overlapping indices along steps in the all-gather ring or utilizing specially designed all-reduce (li2022near). In Section A of the Appendix, we also discuss the communication measured during a live deployment of collaborative learning over the internet using SparseLoCo.
@@
Although our analysis and motivation in the work focuses on aggressively compressing the per-iteration message size we note that in communication settings where efficient all-reduce is already available and preferred, SparseLoCo still provides significant benefit over DiLoCo while incurring no additional memory or compute overhead. Concretely, the aggregation step in Algorithm 1 Line 14 can be performed directly by an all-reduce over a sparse vector. This has two significant benefits over DiLoCo with all-reduce (AR): (1) As observed in Table 9 and Figure 1, the performance when is optimally selected is improved over DiLoCo and (2) the Outer error feedback naturally supports more aggressive quantization than the naive DiLoCo, allowing for 2-bit quantization to be used without an additional accumulator, unlike muloco.
SparseLoCo scales across model sizes, communication intervals, and number of replicas
In Table 4, we evaluate scaling of DiLoCo and SparseLoCo with number of workers using a 512M-parameter model scale and communication interval . Scaling beyond 8 replicas without significant degradation is a known challenge scalingdiloco. We observe that SparseLoCo consistently outperforms DiLoCo with higher number of workers across all settings and across a number of densities, showing that it can help address the challenge of scaling the number of replicas. We also observe that with higher number of workers a lower density can sometimes be supported. In the Appendix 12 we also study the impact of the number of replicas at 178M model size. We also evaluate SparseLoCo in the highest communication intervals (), for this we follow scalingdiloco using an overtraining regime with a doubled token budget, where we see again SparseLoCo is able to achieve competitive performance with DiLoCo while improving communication. Finally, we run a larger scale model of size 2B-parameter model scale with workers and communication interval , where SparseLoCo with density outperforms DiLoCo (Table 5). We can see that the benefits of SparseLoCo are maintained at this scale.
| Method | ARC-Easy | HellaSwag | PIQA |
| AdamW DDP | 44.99% | 36.08% | 65.34% |
| DiLoCo | 44.28% | 34.50% | 64.96% |
@@
| SparseLoCo | 45.24% | 36.49% | 65.23% |
| Method | Density | Loss (R=8) | Loss (R=16) | Loss (R=32) |
| AdamW | 100.00% | 2.69 | 2.69 | 2.69 |
| DiLoCo | 100.00% | 2.81 | 2.87 | 2.93 |
| SparseLoCo | 0.78% | 2.97 | 3.00 | 3.09 |
| 1.56% | 2.89 | 2.92 | 3.00 | |
| 3.12% | 2.83 | 2.86 | 2.92 | |
| 6.25% | 2.79 | 2.82 | 2.88 | |
| 12.50% | 2.78 | 2.80 | 2.91 | |
| 25.00% | 2.78 | 2.84 | 3.02 |
| Method | Val Loss | ARC-Easy | ARC-Challenge | HellaSwag | PIQA | WinoGrande |
| AdamW DDP | 2.34 | 58.42% | 32.00% | 56.87% | 72.25% | 56.59% |
| DiLoCo | 2.37 | 60.48% | 30.55% | 54.95% | 72.85% | 55.56% |
| SparseLoCo | 2.36 | 59.05% | 32.17% | 55.49% | 72.85% | 58.56% |
@@ 4.3 Ablations
A natural way to combine DiLoCo with OuterEF is by adding an error feedback while keeping the Nesterov outer optimizer. This has been attempted by muloco, who showed it can provide benefits for quantization but that performance degrades quickly, despite using EF, with sparsification. This approach requires an additional accumulator. In contrast, our finding is that, in the case of high sparsification, using a global outer momentum can be detrimental to performance. This is illustrated in Table 6, where we equip SparseLoCo’s outer optimizer with Nesterov outer momentum. This significantly degrades performance at high sparsity. We hypothesize that this is due to the conflicting directions of the error feedback and the outer momentum, since the largest components become amplified by the outer momentum but not the error feedback.
| Method | Loss | |
| SparseLoCo | 2.70 | |
3.39 |
@@
We ablate the choice of compared to the alternative Random- (shi2019topk; pmlr-v202-wang23t) in Table 7. We observe that performance is significantly degraded when using random- for the same number of indices selected, emphasizing the importance of this design choice.
@@
As discussed, SparseLoCo benefits from stronger quantization than non-EF methods and supports up to 2-bit quantization and was generally observed to give results very close to full precision. In Table 7, we show the performance at different quantization values, showing that 2-bit quantization can be achieved at almost no performance cost.
Density Random- Loss Top-k Loss 1.56% 3.05 2.74 3.12% 2.98 2.70 6.25% 2.93 2.71 Quant. bits 1 2 3 4 32 Loss 4.79 2.70 2.70 2.70 2.70
@@ 5 Conclusion
We have proposed an algorithm that can blend multi-iteration LLM pre-training methods with sparsification and quantization, enabling aggressive compression of DiLoCo’s pseudo-gradients. Our work establishes that the outer momentum in DiLoCo can be replaced by local momentum accumulators without losing performance. Connecting local momentum with error feedback, we leverage this insight to develop SparseLoCo. Our extensive experiments confirm that SparseLoCo significantly reduces communication while outperforming strong baselines such as DiLoCo and DeMo, placing it on the Pareto frontier of loss versus communication volume. Additionally, our experiments reveal that sparse aggregation may actually be useful for improving the performances, opening the possibility of studying more sophisticated aggregation methods in the pre-training setting.
@@ Appendix A Real-World Deployment for Collaborative Permissionless Distributed Training over the Internet
SparseLoCo has been deployed in a real-world setting and is being used to collaboratively train models up to 8B and 70B with permissionless global participants using an incentive scheme lidin2025incentivizing that rewards participants purely based on analysis of their compressed pseudo-gradients. This was done on top of an existing blockchain. The addition and coordination of peers and their rewards were handled through the blockchain. Communication of pseudo-gradients was routed through globally distributed, S3-compliant object storage—specifically Cloudflare R2—which enabled rapid dissemination of model updates worldwide. This setup allowed updates to be time-stamped and verified as part of the reward mechanism lidin2025incentivizing. Each peer maintained their own storage bucket, posting read credentials to a blockchain so that both other peers and the reward mechanism could access their compressed pseudo-gradients.
SparseLoCo is particularly advantageous in this communication setup, as cloud providers have large bandwidth for peer downloads and are able to rapidly distribute and mirror files across the globe. For upload, a peer only sends their pseudo-gradient through the cloud provider. Therefore, their outbound communication (and required upload bandwidth) is kept low. They then download the pseudo-gradients from the cloud provider, which is able to easily handle the high bandwidth constraints. An example of a practical communication time measured with an 8B model is on average 12 seconds, including sending their compressed pseudo-gradients and downloading other workers’ messages with the test node never exceeding 500 Mb/s. Compared to processing with H200, which takes around 4.5 minutes, leading to minimal wall-clock time degradation despite traffic over the internet. For reference, intellect1, which trained a similarly sized model (10B) with 8-bit DiLoCo, reports a globally distributed all-reduce synchronization time of 8.3 minutes on average for a peak of nodes participating and processing time of 38 minutes. We also performed test measurements of communication time for a 70B model with the same setup as above ( peers), measuring a total communication time of 70 seconds on average, with the test node never reaching more than 500 Mb/s downlink and 110 Mb/s uplink.
70B LLM training We have deployed SparseLoCo on the live system discussed above to train a 72B model, the largest collaborative foundational model training run ever considered. The deployment uses replicas each associated to a peer, inner steps and a global batch size of 8M tokens per inner step. Although the system design allows peers to use any target hardware that can achieve reasonable throughput, the suggested hardware requirements targeted an B200. Preliminary results on several benchmarks after an estimated 120B tokens have yielded bpb of 0.798 and Hellaswag (0-shot) downstream of 71.2% consistent with expectations at this scale, though no existing benchmarks at similar token budgets are available due to the scale of training.
@@ Appendix B Effect of Chunking, DCT, and inner steps
A recently introduced method peng2024decoupled considered the single-step setting with error feedback, using a compression function that first applies a discrete cosine transform (DCT) on tensor chunks and then selects the values in the DCT domain. It further employed sign descent on the final aggregated update kunstner2023noise. Without the DCT transform, this approach can be seen as a special case of SparseLoCo when , the inner optimizer is plain SGD, and the outer optimizer utilizes sign descent. Since the effect of the DCT transform, designed for data with sequential structure where the order of elements matters, is not well understood in this context, and given the additional uncertainty about the role of chunking, in this section we disentangle the contributions of both for DeMo, as well as in the multi-step () setting. The ablations of these three factors, evaluated purely in terms of loss, are presented in Table 8. Here, the EF baseline is a simplified DeMo that applies selection globally to the entire tensor (rather than within chunks) while still utilizing sign descent.
@@
Notably, DeMo does not have a natural way to incorporate adaptive optimization, and in practice, the sign descent is used to approximate the benefits of the Adam optimizer peng2024decoupled. A significant advantage of SparseLoCo is that operating on the pseudo-gradients allows easy integration of adaptive optimizers like Adam in the inner loop.
@@ Appendix C Overtraining Regime with Large Communication Interval
Following scalingdiloco, we put SparseLoCo to the test in an overtraining regime by doubling the token budget to 20B and using a larger communication interval of . Our observations are consistent with the trends in Figure 1, and SparseLoCo outperforms DiLoCo at this setting (Table 9).
@@
Appendix D Streaming SparseLoCo
In this section, we verify that Streaming DiLoCo streamingdiloco and SparseLoCo can be combined. Streaming DiLoCo is an orthogonal direction to SparseLoCo for reducing peak communication volume by hiding it. SparseLoCo reduces the absolute number of bits communicated per step through compression, thus indirectly reducing peak communication. Streaming DiLoCo directly reduces peak bandwidth by only communicating subsets of the model’s parameters at a time but does not reduce the absolute number of bits communicated.
Table 10 reports results for combining SparseLoCo with Streaming DiLoCo to reduce peak communication volume. We train 18-layer 1B-parameter transformers (hidden dimension ) in this ablation. The model is partitioned into three even subsets of hidden layers, with the first and third subsets containing the embedding and unembedding layers, respectively. We train the M parameter model for a chinchilla-optimal B tokens hoffmann2022training. We use a communication interval of for the full model (Streaming communicates every steps) and workers. We observe that both models reach the same final validation loss (it differed only in the 4th decimal), while Streaming SparseLoCo reduces peak communication volume by a factor of .
| Method | Density | Comm. Volume/Step | Peak Comm. Volume | Loss |
| SparseLoCo | 3.125% | 35.03 MB | 35.03 MB | 2.51 |
| Streaming SparseLoCo | 3.125% | 35.03 MB | 11.68 MB | 2.51 |
Appendix E Freezing Error Feedback
We apply a short error feedback (OuterEF) freeze at the beginning of training: for the first few outer steps, the error feedback is not utilized. Concretely, during the freeze we don’t use nor accumulate in the EF buffer. We find that freezing the OuterEF for the first few outer steps slightly improves training stability and overall performance (see Table 11).
| EF Freeze | Loss |
| 0% | 2.704 |
| 5% | 2.699 |
Appendix F Scaling Replicas Across Different Densities and Communication Intervals
We compare DiLoCo and SparseLoCo while varying the number of workers , communication intervals using model sizes 178M and 512M, and report the final validation loss in Tables 12, 13, and 4. We observe that SparseLoCo outperforms DiLoCo with higher number of parallel workers.
H=15
Method
Density
Loss
R=8
R=32
AdamW
100.00%
2.91
2.91
DiLoCo
100.00%
2.99
3.10
SparseLoCo
0.78%
2.96
3.02
1.56%
2.93
3.00
3.12%
2.91
2.99
6.25%
2.94
3.00
12.50%
2.96
3.04
25.00%
2.96
3.14
50.00%
3.03
3.29
H=50
Method
Density
Loss
R=8
R=32
AdamW
100.00%
2.91
2.91
DiLoCo
100.00%
3.05
3.20
SparseLoCo
0.78%
3.09
3.13
1.56%
3.03
3.07
3.12%
2.99
3.09
6.25%
2.98
3.09
12.50%
3.00
3.14
25.00%
3.04
3.25
50.00%
3.12
3.42
H=100
Method
Density
Loss
R=8
R=32
AdamW
100.00%
2.91
2.91
DiLoCo
100.00%
3.12
3.29
SparseLoCo
0.78%
3.20
3.29
1.56%
3.12
3.26
3.12%
3.05
3.19
6.25%
3.03
3.17
12.50%
3.03
3.21
25.00%
3.07
3.32
50.00%
3.17
3.48
| Method | Density | Loss (R=8) | Loss (R=16) | Loss (R=32) |
| AdamW | 100.00% | 2.69 | 2.69 | 2.69 |
| H=15 | ||||
| DiLoCo | 100.00% | 2.76 | 2.77 | 2.82 |
| SparseLoCo | 0.78% | 2.79 | 2.81 | 2.84 |
| 1.56% | 2.74 | 2.76 | 2.79 | |
| 3.12% | 2.70 | 2.74 | 2.77 | |
| 6.25% | 2.71 | 2.76 | 2.78 | |
| 12.50% | 2.76 | 2.78 | 2.82 | |
| 25.00% | 2.77 | 2.78 | 2.93 | |
| H=100 | ||||
| DiLoCo | 100.00% | 2.87 | 2.94 | 3.05 |
| SparseLoCo | 0.78% | 3.09 | 3.14 | 3.29 |
| 1.56% | 3.02 | 3.06 | 3.21 | |
| 3.12% | 2.96 | 3.03 | 3.12 | |
| 6.25% | 2.94 | 2.97 | 3.03 | |
| 12.50% | 2.85 | 2.88 | 3.02 | |
| 25.00% | 2.82 | 2.89 | 3.11 | |
Appendix G Compression of indices in
In methods, the indices of the selected values need to be transmitted alongside the values. When values are aggressively quantized (as in SparseLoCo), this index-transmission overhead becomes significant. In SparseLoCo, we utilize chunk sizes of , so, naively, we can transmit indices in bits per transmitted value. However, with 2-bit quantization, this overhead becomes significant, motivating further index compression. Assuming a chunk size of and selection, we observe that the information-theoretic limit is bits. For practical cases considered in this work ( and ), this corresponds to 8.3, 6.3, and 5.3 bits per transmitted value, respectively. In practice, we designed a custom compression algorithm based on sub-chunking and coding that achieves 8.9, 6.6, and 5.6 bits per value for these cases.
Appendix H Hyperparameter Selection Details
Table 15 reports the hyperparameter search spaces for the 512M model size. We tune all methods at communication interval and reuse the best configurations for other settings; General and model architecture settings are fixed across all runs unless stated otherwise. In Table 14, we provide architectural details for 178M and 2B model sizes. For 178M, we reduce the batch size to 32 leading to an effective batch size of , and repeating the hyper-parameter sweeps as Table 15, we observe the same optimal settings. For the 2B model size, we increase warmup to and perform a small sweep of learning-rates lower than the optimal setting of 178M and 500M models. Specifically, for DiLoCo we search and , finding , optimal; for SparseLoCo we search and , finding , optimal. For number of workers experiments, we ensure the same effective batch size used for by scaling the batch size accordingly.
| Parameter | Value |
| Total Parameters | 177,622,016 |
| Number of Layers | 9 |
| Hidden Size | 1,024 |
| Intermediate Size | 2,688 |
| Attention Heads | 8 |
| Vocabulary Size | 32,000 |
| FFN Activation | SwiGLU |
| Parameter | Value |
| Total Parameters | 1,972,759,040 |
| Number of Layers | 24 |
| Hidden Size | 2,560 |
| Intermediate Size | 7,680 |
| Attention Heads | 20 |
| Key-Value Heads | 5 |
| Vocabulary Size | 32,000 |
| FFN Activation | SwiGLU |